# Localizing Dynamically-Formed Black Hole Binaries in Milky Way Globular Clusters with LISA

### v1

@@ Abstract
The dynamical formation of binary black holes (BBHs) in globular clusters (GCs) may contribute significantly to the observed gravitational wave (GW) merger rate. Furthermore, LISA may detect many BBH sources from GCs at mHz frequencies, enabling the characterization of such systems within the Milky Way and nearby Universe. In this work, we use Monte Carlo N-body simulations to construct a realistic sample of Galactic clusters, thus estimating the population, detectability, and parameter measurement accuracy of BBHs hosted within them. In particular, we show that the GW signal from $0.7\pm 0.7$, $2.0\pm 1.7$, $3.6\pm 2.3$, $13.4\pm 4.7$ BBHs in Milky Way GCs can exceed the signal-to-noise ratio threshold of $\rm SNR=30$, 5, 3, and 1 for a 10-year LISA observation, with $\sim 50\%$ of detectable sources exhibiting high eccentricities ($e\gtrsim 0.9$). Moreover, the Fisher matrix and Bayesian analyses of the GW signals indicate these systems typically feature highly-resolved orbital frequencies ($\delta f_{\rm orb}/f_{\rm orb}\sim 10^{-7}-10^{-5}$) and eccentricities ($\delta e/e\sim 10^{-3}-0.1$), as well as a measurable total mass when SNR exceeds $\sim 20$. Notably, BBHs with $\rm SNR\gtrsim 20$ can be confidently localized to specific GCs within the Milky Way, with an angular resolution of $\sim 10-100$ arc minutes. The detection and localization of even a single BBH in a Galactic GC would allow accurate tracking of its long-term orbital evolution, enable a direct test of the role of GCs in BBH formation, and provide a unique probe into the evolutionary history of Galactic clusters.
@@ 2 Creating a mock Galactic sample
For this study, we aim to predict which specific Galactic GCs are most likely to host resolvable LISA sources at present. In this case, we must identify a single “best-fit” model from the CMC Cluster Catalog for each observed Milky Way GC. The CMC Catalog models can be sorted into a three-dimensional grid in $M_{\rm cl}-Z-R_{\rm gc}$ space, where $M_{\rm cl}$ is the total cluster mass at present (set by our choice of initial $N=[2,4,8,16,32]\times 10^{5}$ stars), $Z=[0.01,0.1,1]\times Z_{\odot}$ is the cluster metallicity, and $R_{\rm gc}=[2,8,20]\,$kpc is the Galactocentric position in the Galaxy. For a given observed GC, we identify the appropriate $M_{\rm cl}-Z-R_{\rm gc}$ model bin based on that GC’s observed mass and metallicity values (taken from Harris, 2010). Each distinct $M_{\rm cl}-Z-R_{\rm gc}$ bin contains 4 models of varying initial virial radius, which determines the clusters’ final core and half-light radius. To find the single best-fit model, we identify the model in the appropriate $M_{\rm cl}-Z-R_{\rm gc}$ bin which has core radius at $t=12\,$Gyr closest to the observed core radius value (Harris, 2010). [^1] We also tried fitting using half-light radius in place of core radius, and found no major differences in our results. Once the best fit model is identified, we select a cluster age by randomly sampling from that model’s available time snapshot outputs in the range $8-13.7\,$Gyr (typically this consists of a sample of 10-50 snapshots per model). This range is intended to reflect the uncertainty in cluster ages at present. Once a cluster snapshot is selected, we then identify the number and properties of all BBHs present in the best-fit model at that time. By repeating these steps for each observed GC, we build a sample of the full population of BBHs present in each of the Milky Way’s GCs at present. We then repeat this full procedure ten times to assemble ten separate Galactic realizations (the random time snapshot draw enables us to create distinct ten distinct samples).
@@ 3.2 The Location of Gravitational Wave Sources
The results are summarized in Figure 2. Specifically, Left Panel depicts the ratio of the clusters’ core radius to the half-light radius, $r_{c}/r_{h}$ (y-axis), versus their distance, $R$, from the detector (x-axis); Right Panel shows the estimated total mass of each cluster, $M_{\rm GC}$, against their distance $R$ (x-axis). The value of $r_{c}/r_{h}$ and $R$ are taken from Harris (2010), and the value of $M_{\rm GC}$ is adopted from Baumgardt & Hilker (2018). In Figure 2, the size of solid dots represents the expected probability for a cluster to host a BBH system with $\rm SNR>1$:
@@ 3.3 Astrophysical Implication
After generating the GW signal as described by Equation (9), we compute the partial derivatives of the waveform with respect to each parameter (see Equation (7)). For example, to calculate $\partial h/\partial M$, we vary the total mass $M\rightarrow M^{\prime}=M+dM$, and generate a new waveform $h^{\prime}(t)=h^{\prime}(t;\boldsymbol{\lambda}^{\prime})$, where $\boldsymbol{\lambda}^{\prime}=\left\{f_{\rm orb},1-e,M+dM,q,\cos\Theta,\Phi,\cos\theta,\phi,R,\psi\right\}$. The partial derivative is then approximated as $\partial h/\partial M\approx[h^{\prime}(t)-h(t)]/\Delta M$. Note that each partial derivative is a time series representing the difference in the waveform caused by slightly varying one of the parameters around the central value. Also, in the numerical computation, we choose the parameter variation $d\lambda_{i}$ such that $\left\langle\left.dh\right\rvert\,dh\right\rangle\sim 10^{-3}\left\langle\left.h\right\rvert\,h\right\rangle$, ensuring the change in the waveform is small when computing the numerical derivative.
@@
We highlight that, the estimation shown in Figure 3 is agnostic to different BBH formation channels. In other words, for any potential BBH population in the galaxy, not necessarily from the GCs, their orbital parameter $(a,1-e)$ can be over-plotted on the figure to estimate the parameter measurement errors in a similar way. Moreover, Figure 3 indicates that our analysis of the parameter measurement accuracy is robust to variations in simulation results. For example, in Panels $a$ and $b$, regions with $\rm SNR>1$ (to the left of the dashed line at $\rm SNR=1$) show $\delta f_{\rm orb}/f_{\rm orb}\lesssim 10^{-5}$ and $\delta e/e\lesssim 10\%$. This suggests that, in realistic observations, BBH systems detected in the Milky Way GCs can typically achieve high accuracy measurements of $f_{\rm orb}$ and $e$. Furthermore, as shown by Panel c, BBHs with $\rm SNR>20$ can have a mass measurement accuracy of $\delta M/M\lesssim 10\%$. However, marginally detectable sources may have poorly constrained total mass (blue regions near the dashed line of $\rm SNR=5$). Similarly, Panel d shows that BBHs with $\rm SNR>20$ can have a sky localization accuracy of $2a_{s}\lesssim 300$ arcmins (yellow regions), which indicates that high-SNR BBHs can generally be localized with an angular resolution of a few degrees in the Milky Way.
@@
Figure 5: Parameter Measurement error of an example BBH system in Milky Way Globular Clusters, estimated using Bayesian analysis. Here we consider the GW signal from a BBH system in the simulation, with intrinsic parameters of $m_{1}=23.3$ ${}_{\odot},$ $m_{2}=18.6$ ${}_{\odot},$ $a=0.06332$ au ($f_{\rm orb}=1.28771635\times 10^{-5}$ Hz), $e=0.897,$ $R=4.9$ kpc, assuming $\Phi=\Theta=\theta=\phi=\psi=\pi/4$ for simplicity. The color maps show the posterior distributions of the binary’s observed parameters (1-$\sigma$ to 5-$\sigma$ level, from red to blue), assuming a uniform a priori distribution and stationary Gaussian LISA noise. Note that this system has a SNR of $\sim 75$ for 5-year observation. For demonstration purposes, we use the cross in each panel to indicate the system’s intrinsic parameters. Furthermore, to avoid high computational cost, we do not include a full Bayesian analysis of all the 10 parameters, but only treat two of them as random variables in each panel (see the x and y axes) and fix the value of others.
@@
Furthermore, we highlight that, LISA can accurately localize most BBHs with $\rm SNR>20$ in Milky Way GCs (see the red and orange dots in Bottom Panel). In particular, these GW sources have a sky localization accuracy of $\sim 10-100$ arcmins (i.e., less than a few degrees), which is comparable to the typical tidal radius of Milky Way GCs ($\sim 100-200$ arcmins, see, e.g., Harris (2010)). As a result, LISA can localize BBH sources in specific Milky Way GCs, providing unique information about the Milky Way’s compact binary formation and the dynamical environment of GCs. Additionally, while the distance of BBHs is less precisely constrained compared to their sky location, most systems with $\rm SNR>20$ have a distance measurement error of $\Delta R\sim 1-100$ kpc in the Milky Way. Given the significant distance to the nearby galaxies (e.g., the Andromeda Galaxy is located at $\sim 780$ kpc), we expect these BBHs to be distinguishable from extragalactic GW sources.
@@
We note that, although Fisher matrix analysis has been widely used to estimate the parameter measurement error for LISA sources, this method can sometimes yield inaccurate results, particularly for degenerate parameters or parameters with weak influences on the shape of GW signal (see, e.g., Vallisneri, 2008; Toubiana et al., 2020). Consequently, the parameter measurement accuracy presented in Figures  3 and 4 should be interpreted as a heuristic estimation, (especially for the mass parameters $M$ and $q$; see, e.g., Xuan et al., 2024b, for more details). Nevertheless, these estimates provide compelling evidence that BBHs in Milky Way GCs are promising targets for inferring their astrophysical properties.
To further validate the Fisher matrix results, we performed a Bayesian analysis (see, e.g., Finn, 1992; Cutler & Flanagan, 1994; Christensen & Meyer, 1998; Christensen et al., 2004), as shown in Figure 5. In particular, we choose an example BBH system from the simulation results (intrinsic parameters $m_{1}=23.3$ ${}_{\odot},$ $m_{2}=18.6$ ${}_{\odot},$ $a=0.06332$ au, $e=0.897,$ $R=4.9$ kpc), and inject its GW signal $s_{0}(t;\boldsymbol{\lambda})$ into a simulated stationary Gaussian LISA noise $n(t)$. We then vary the parameters slightly around their intrinsic values of $\boldsymbol{\lambda}$, generate different GW templates $s(t;\boldsymbol{\lambda}^{\prime})$, and compute their inner product with the mock signal $z(t)=s_{0}(t;\boldsymbol{\lambda})+n(t)$ (see Equation (6)). The posterior distribution of the example binary’s observed parameters is then computed using the aforementioned inner products $\langle s\mid z\rangle$, following equations 1-6 in Ref. Christensen et al. (2004). Due to the high computational cost, our comparison does not include a full Bayesian analysis of all the 10 parameters but only varies two of them in each panel (see the x and y axes) and fixes the others. We compute the results for five representative parameters, $f_{\rm orb}$, $e$, $M$, $R$, and $\phi$ (which is one of the sky location angles, representing the source’s sky localization accuracy).
As shown by Figure 5, despite the degeneracies and bias of some parameters, Bayesian analysis yields their measurement error (as indicated by the size of the 3-$\sigma$ level contours in Figure 4) comparable with the Fisher matrix estimation. Particularly, the Fisher matrix estimation of this example BBH system is represented by the red triangle to the bottom-left of each panel in Figure 4, with its parameter measurement error $\delta f_{\rm orb}\sim 2.1\times 10^{-12}$ Hz, $\delta e\sim 1.4\times 10^{-3},$ $\delta M\sim 1.4$ ${}_{\odot},$ $\delta R\sim 7.8$ kpc, and sky localization accuracy $2a_{s}\sim 39$ arcmin0.65similar-toabsent0.65$\sim$ 0.65∼ 0.65 degrees. These values are consistent with the size of 3-$\sigma$ contours in Bayesian analysis (within one order of magnitude, see Figure 5), which partly justifies the Fisher matrix results of Figures 3 and 4.

### v2

@@ Abstract
The dynamical formation of binary black holes (BBHs) in globular clusters (GCs) may contribute significantly to the observed gravitational wave (GW) merger rate. Furthermore, LISA may detect many BBH sources from GCs at mHz frequencies, enabling the characterization of such systems within the Milky Way and nearby Universe. In this work, we use Monte Carlo N-body simulations to construct a realistic sample of Galactic clusters, thus estimating the population, detectability, and parameter measurement accuracy of BBHs hosted within them. In particular, we show that the GW signal from $0.7\pm 0.7$, $2.0\pm 1.7$, $3.6\pm 2.3$, $13.4\pm 4.7$ BBHs in Milky Way GCs can exceed the signal-to-noise ratio threshold of $\rm SNR=30$, 5, 3, and 1 for a 10-year LISA observation, with $\sim 50\%$ of detectable sources exhibiting high eccentricities ($e\gtrsim 0.9$). Moreover, the Fisher matrix and Bayesian analyses of the GW signals indicate these systems typically feature highly-resolved orbital frequencies ($\delta f_{\rm orb}/f_{\rm orb}\sim 10^{-7}-10^{-5}$) and eccentricities ($\delta e/e\sim 10^{-3}-0.1$), as well as a measurable total mass when SNR exceeds $\sim 20$. Notably, we show that high-SNR BBHs can be confidently localized to specific Milky Way GCs with a sky localization accuracy of $\delta\Omega\sim 1$ deg2, and address the large uncertainties in their distance measurement ($\delta R\sim 0.3-20$ kpc). The detection and localization of even a single BBH in a Galactic GC would allow accurate tracking of its long-term orbital evolution, enable a direct test of the role of GCs in BBH formation, and provide a unique probe into the evolutionary history of Galactic clusters.
@@ 2 Creating a mock Galactic sample
For this study, we aim to predict which specific Galactic GCs are most likely to host resolvable LISA sources at present. In this case, we must identify a single “best-fit” model from the CMC Cluster Catalog for each observed Milky Way GC. The CMC Catalog models can be sorted into a three-dimensional grid in $M_{\rm cl}-Z-R_{\rm gc}$ space, where $M_{\rm cl}$ is the total cluster mass at present (set by our choice of initial $N=[2,4,8,16,32]\times 10^{5}$ stars and evolved to the current value), $Z=[0.01,0.1,1]\times Z_{\odot}$ is the current cluster metallicity, and $R_{\rm gc}=[2,8,20]\,$kpc is the current Galactocentric position. For a given observed GC, we identify the appropriate $M_{\rm cl}-Z-R_{\rm gc}$ model bin based on that GC’s observed mass and metallicity values (taken from Harris, 2010). Each distinct $M_{\rm cl}-Z-R_{\rm gc}$ bin contains 4 models of varying initial virial radius, which determines the clusters’ final core and half-light radius. To find the single best-fit model, we identify the model in the appropriate $M_{\rm cl}-Z-R_{\rm gc}$ bin which has core radius at $t=12\,$Gyr closest to the observed core radius value (Harris, 2010). [^1] We also tried fitting using half-light radius in place of core radius, and found no major differences in our results. Once the best fit model is identified, we select a cluster age by randomly sampling from that model’s available time snapshot outputs in the range $8-13.7\,$Gyr (typically this consists of a sample of 10-50 snapshots per model). This range is intended to reflect the uncertainty in cluster ages at present. Once a cluster snapshot is selected, we then identify the number and properties of all BBHs present in the best-fit model at that time. By repeating these steps for each observed GC, we build a sample of the full population of BBHs present in each of the Milky Way’s GCs at present. We then repeat this full procedure ten times to assemble ten separate Galactic realizations (the random time snapshot draw enables us to create distinct ten distinct samples).
@@ 3.2 The Location of Gravitational Wave Sources
The results are summarized in Figure 2. Specifically, Left Panel depicts the ratio of the clusters’ core radius to the half-light radius, $r_{c}/r_{h}$ (y-axis), versus their distance, $R$, from the detector (x-axis); Right Panel shows the estimated total mass of each cluster, $M_{\rm GC}$, against their distance $R$ (x-axis) (see Harris (2010)). In Figure 2, the size of solid dots represents the expected probability for a cluster to host a BBH system with $\rm SNR>1$:
@@ 3.3 Astrophysical Implication
After generating the GW signal as described by Equation (9), we compute the partial derivatives of the waveform with respect to each parameter (see Equation (7)). For example, to calculate $\partial h/\partial M$, we vary the total mass $M\rightarrow M^{\prime}=M+\Delta M$, and generate a new waveform $h^{\prime}(t)=h^{\prime}(t;\boldsymbol{\lambda}^{\prime})$, where $\boldsymbol{\lambda}^{\prime}=\left\{f_{\rm orb},1-e,M+\Delta M,q,\cos\Theta,\Phi,\cos\theta,\phi,R,\psi\right\}$. The partial derivative is then approximated as $\partial h/\partial M\approx[h^{\prime}(t)-h(t)]/\Delta M$. Note that each partial derivative is a time series representing the difference in the waveform caused by slightly varying one of the parameters around the central value. Also, in the numerical computation, we choose the parameter variation $d\lambda_{i}$ such that $\left\langle\left.dh\right\rvert\,dh\right\rangle\sim 10^{-3}\left\langle\left.h\right\rvert\,h\right\rangle$, ensuring the change in the waveform is small when computing the numerical derivative.
@@
We highlight that the estimation shown in Figure 3 is agnostic to different BBH formation channels. In other words, for any potential BBH population in the galaxy, not necessarily from the GCs, their orbital parameter $(a,1-e)$ can be over-plotted on the figure to estimate the parameter measurement errors in a similar way. Moreover, Figure 3 indicates that our analysis of the parameter measurement accuracy is robust to variations in simulation results. For example, in Panels $a$ and $b$, regions with $\rm SNR>1$ (to the left of the dashed line at $\rm SNR=1$) show $\delta f_{\rm orb}/f_{\rm orb}\lesssim 10^{-5}$ and $\delta e/e\lesssim 10\%$. This suggests that, in realistic observations, BBH systems detected in the Milky Way GCs can typically achieve high accuracy measurements of $f_{\rm orb}$ and $e$. Furthermore, as shown by Panel c, BBHs with $\rm SNR>20$ can have a mass measurement accuracy of $\delta M/M\lesssim 10\%$. However, marginally detectable sources may have poorly constrained total mass (blue regions near the dashed line of $\rm SNR=5$). Similarly, Panel d shows that BBHs with $\rm SNR>20$ can have a sky localization accuracy of $2a_{s}\lesssim 300$ arcmins (yellow regions), which indicates that high-SNR BBHs can generally be localized with an angular resolution of a few degrees in the Milky Way.
@@
Figure 5: Parameter Measurement error of an example BBH system in Milky Way Globular Clusters, estimated using Bayesian analysis (for 5-yr observation). Here we consider a BBH system with $m_{1}=23.3$ ${}_{\odot},$ $m_{2}=18.6$ ${}_{\odot},$ $f_{\rm orb}=1.2877164\times 10^{-5}$ Hz, $e=0.897,$ $R=4.9$ kpc, $\Phi=158.7^{\circ},\,\Theta=141.4^{\circ},\,\phi=219.9^{\circ},\,\theta=125.2^{\circ},\,\psi=153.9^{\circ}$. (Note that this system has $\rm SNR\sim 77$, and we keep 8 significant digits of $f_{\rm orb}$ to highlight its realistic measurement accuracy.) The color maps show the marginal posterior density of the binary’s observed parameters, assuming a flat, 10-dimensional hyperrectangular prior and stationary Gaussian LISA noise. We use the star in each panel to indicate the system’s intrinsic parameters, and plot 68%, 95%, and 99.7% credible regions for the marginal posterior densities (i.e, $1-3\,\sigma$ level sets defined analogously to a bivariate Gaussian density). To avoid clutter, we plot the results for the binary’s intrinsic parameters $f_{\rm orb},\,e,\,M$; distance $R$; and sky location $\phi,\,\theta$, which are mostly relevant to its astrophysical inference. Additionally, the peak of posterior density exhibits a small offset from some parameters’ intrinsic value (typically within $1\sigma$ level), which is caused by the random fluctuation of time-domain LISA noise and the limit of numerical accuracy for highly eccentric binary waveforms.
@@
Furthermore, we highlight that, LISA can accurately localize most BBHs with $\rm SNR>20$ in Milky Way GCs (see the red and orange dots in Bottom Panel). In particular, these GW sources have a sky localization accuracy of $\sim 10-100$ arcmins (i.e., less than a few degrees), which is comparable to the typical diameter of Milky Way GCs ($\sim 10-120$ arcmins  [^8] Note that detectable BBHs are more likely to be hosted by nearby, massive GCs (see Section 3.2); as a result, their host clusters tend to have larger angular sizes than the general Milky Way GC population. For BBHs with $\rm SNR>1$, the average host GC diameter is $\sim 38$ arcmins (see Figure 6 in the Appendix). ). Additionally, while the distance of BBHs is less precisely constrained compared to their sky location, most systems with $\rm SNR>20$ have a distance measurement error of $\delta R\sim 1-100$ kpc in the Milky Way. Given the significant distance to the nearby galaxies (e.g., the Andromeda Galaxy is located at $\sim 780$ kpc), we expect these BBHs to be distinguishable from extragalactic GW sources.
3.4 Further validation of BBH localization using Bayesian Analysis
As shown by Section 3.3, LISA has the potential to confidently localize a BBH system in specific Milky Way GCs, which could provide unique information about the Milky Way’s compact binary formation and the dynamical environment of GCs. In this section, we further verify this point using a Bayesian analysis and address the uncertainty introduced by the BBHs’ distance measurement.
@@ 3.4 Further validation of BBH localization using Bayesian Analysis
First, we note that, although the Fisher matrix analysis adopted in Section 3.3 has been widely used to estimate the parameter measurement error for LISA sources, this method can sometimes yield inaccurate results, particularly for degenerate parameters or parameters with weak influences on the shape of GW signal (see, e.g., Vallisneri, 2008; Toubiana et al., 2020). Consequently, the parameter measurement accuracy presented in Figures  3 and 4 should be interpreted as a heuristic estimation. Nevertheless, these estimates provide compelling evidence that BBHs in Milky Way GCs are promising targets for inferring their astrophysical properties.
To further validate the Fisher matrix results, we performed a full 10-dimensional Bayesian analysis for individual systems (see, e.g., Finn, 1992; Cutler & Flanagan, 1994; Christensen & Meyer, 1998, 2001), as shown in Figure 5 and Table 1. Particularly, in Figure 5, we choose an example BBH system from the simulation results ($m_{1}=23.3$ ${}_{\odot},$ $m_{2}=18.6$ ${}_{\odot},$ $a=0.06332$ au, $e=0.897,$ $R=4.9$ kpc), and inject its GW signal $s_{0}(t;\boldsymbol{\lambda})$ [^9] Note that this system has $\rm SNR\sim 77$ for 5-yr observation, and the detector response function is included in $s_{0}(t)$, see Section 3.3. into a simulated stationary Gaussian LISA noise $n(t)$, which yields the mock LISA signal:
$z(t)=s_{0}(t;\boldsymbol{\lambda})+n(t)\ .$ (10)
We then adopted Monte-Carlo sampling, generating GW templates $s(t;\boldsymbol{\lambda}^{\prime})$ with parameters slightly different from their intrinsic values of $\boldsymbol{\lambda}$, and computing their inner product with the mock signal (see Equation (6)):
$\left\langle z\mid s\right\rangle=2\int_{0}^{\infty}\frac{\tilde{z}(f)\tilde{s}^{*}(f)+\tilde{z}^{*}(f)\tilde{s}(f)}{S_{\mathrm{n}}(f)}\mathrm{d}f\ .$ (11)
Assuming a uniform hyperrectangular prior, the posterior probability density of the observed parameters can be computed using the aforementioned inner products (see, e.g., Eqs 1-6 in Christensen & Meyer (2001)):
$p(\boldsymbol{\lambda}^{\prime}\mid z)\propto p(z\mid\boldsymbol{\lambda}^{\prime})=K\exp[2\langle z\mid s(\boldsymbol{\lambda}^{\prime})\rangle-\langle s(\boldsymbol{\lambda}^{\prime})\mid s(\boldsymbol{\lambda}^{\prime})\rangle]\ ,$ (12)
where K is a constant.
[…]
Using the aforementioned method, we showcase three representative systems from the mock Milky Way GC catalog, with $e=0.1671-0.8969$ (see Table 1 in the appendix). To better describe the sources’ localization accuracy, we convert the posterior of their sky location angles, $(\phi,\,\theta)$, into the solid angle of the corresponding sky area $\delta\Omega_{\rm Bayesian}$ ($68\%$ credible region). We then compare $\delta\Omega_{\rm Bayesian}$ with the area of sky error ellipsoid from Fisher matrix analysis, $\delta\Omega_{\rm Fisher}$, and the size of their host GCs. As shown in Table 1, the Fisher matrix and Bayesian analysis generally yield consistent results in the sources’ sky localization accuracy, although the Fisher results are systematically smaller by a factor of $\sim 2$ in angular size estimation (or $\sim 5$ in sky area). Furthermore, for high-SNR BBHs (see, e.g., the three systems with $\rm SNR=35-94$ in Table 1), their sky area accuracy from both Fisher matrix and Bayesian analysis, $\sim 0.40,1.00,1.85$ deg2, are comparable with (or even smaller than) the size of their host clusters, $\sim 2.84,2.13,0.52$ deg2.101010However, marginally detectable BBHs ($\rm SNR\sim 8$) can have poor sky localization. For example, the system in Figure 7 has a sky localization error of $\sim$ 233  deg2 in Bayesian analysis, far exceeding the angular size of its host cluster NGC7078 ($\sim 0.39$ deg2). These results further confirm that BBHs in the Milky Way can be localized to the sky area of specific host GCs, especially for significant detection cases.
Notably, Figures 4-5 and Table 1 indicate a distance measurement error of $\delta R\sim 1-100$ kpc for the Milky Way BBH sources, which is much larger than the tidal diameter of globular clusters (typically on the order of a few tens of pc). Therefore, LISA is unlikely to directly associate the spatial position of a BBH system with its host GC because of the large uncertainty in the radial distance measurement. However, when determining the host environment of BBH sources, this uncertainty can be resolved, mostly because stellar density in the Milky Way halo is much lower than that of GCs. In other words, the total number of stars in a BBH-hosting GC ($\sim 10^{5}-10^{6}$, see Figure 2) is much larger than the total number of halo stars in other parts of the error volume [^11] Note that most GCs reside in the Galatic halo (see,e.g., Harris, 2010)., which makes it highly possible that a detected source is hosted in the cluster instead of other regions of the local universe.
For example, assuming a halo stellar mass density of $\rho_{\rm halo}\sim 10^{-4}\rm M_{\odot}/pc^{3}$ (Deason et al., 2019), the total halo stellar mass in the error volume $V_{\rm err}$ of GW detection can be estimated as:
$M_{\rm halostars}=\rho_{\rm halo}\times V_{\rm err}\sim\rho_{\rm halo}R^{2}\delta R\delta\Omega_{\rm Bayesian}\ ,$ (13)
which yields $M_{\rm halostars}\sim 51\rm M_{\odot},1.5\times 10^{3}\rm M_{\odot},6.5\times 10^{4}\rm M_{\odot}$ for the first three confidently-localized systems in Table 1. These values are much smaller than the total stellar mass of their host GCs, which is estimated as $M_{\rm GC}\sim 7\times 10^{5}\rm M_{\odot},1.5\times 10^{6}\rm M_{\odot},2\times 10^{5}\rm M_{\odot}$ for NGC104, NGC5139, and NGC6341, respectively (see, e.g., Marks & Kroupa, 2010).
We note that the relative contributions of GCs and the Galactic field to the total BBH merger rate remain uncertain (see Mandel & Broekgaarden, 2022, for a review). However, recent studies suggest both are likely contributing at a comparable rate (see, e.g., Zevin et al., 2021; Fishbach & Fragione, 2023). Furthermore, even if the galactic field has a larger contribution to the total merger rate, considering the significant difference in total stellar mass between the field ($\sim 10^{10}\rm M_{\odot}$ for the Milky Way; Cautun et al., 2020) and GCs ($\sim 10^{7}\rm M_{\odot}$ for the Milky Way; Harris, 2010), the BBH merger rate per unit stellar mass in GCs is likely comparable to, or higher than in the field. Thus, we assume that all stellar mass (in the Galactic halo and in globular clusters) has an equal probability of hosting a BBH source. Under this assumption, the probability that a given BBH resides in a GC can be estimated as:
$P_{\rm incluster}\sim\frac{M_{\rm GC}}{M_{\rm GC}+M_{\rm halostars}}\ .$ (14)
Therefore, once a source is accurately located in the sky area of a GC, the probability for the system to be a GC BBH source can be significantly boosted, especially because a globular cluster contributes to the majority of stellar mass in the error volume. Additionally, the highly eccentric formation nature of BBHs in GC (see, e.g., Section 3.1) can also help with distinguishing them from other BBHs from isolated evolution channels in the galactic field, which are expected to have negligible eccentricities (e.g., Belczynski et al., 2016; Stevenson et al., 2017; Eldridge et al., 2019).  [^12] However, the localization method discussed earlier in this section is agnostic to the expected eccentricity distributions of different BBH formation channels. For example, if a population of eccentric BBHs is accurately localized toward the Galactic field rather than any GCs, this method can also distinguish it from the in-cluster BBH population.